. If \\(θ(x,y,z)=3x^2 y-y^3 z^2\\),find \\( ˆ‡Î¸\\) (grad θ) at the point (1, -2, -1).
1
Expert's answer
2020-11-17T16:19:45-0500

"f(x, y, z)=3x^2y -y^3z^2, M(1, -2, -1)"


"\dfrac{\partial f}{\partial x}=6xy,"


"\dfrac{\partial f}{\partial y}=3x^2-3y^2z^2,"

"\dfrac{\partial f}{\partial z}=-2y^3z"

"\dfrac{\partial f(M)}{\partial x}=6(1)(-2)=-12,"

"\dfrac{\partial f(M)}{\partial y}=3(1)^2-3(-2)^2(-1)^2=-9,"

"\dfrac{\partial f(M)}{\partial z}=-2(-2)^3(-1)=-16"


"\text{grad}f(M)=-12\vec{i}-9\vec{j}-16\vec{k}"